If a motorcycle is traveling at 45mph and begins to slide 18 ft before hitting a parked vehicle, how far will the occupants travel before they land once ejected from the motorcycle?
1
Expert's answer
2019-09-02T09:15:37-0400

Conversions:

"1mph=0.44704\frac{meters}{second}"

"1feet=0.3048 meters"

Given that,

Speed

"v=45mph=45\times 0.44704=20.1168\frac{m}{s}"

Skids distance

"d=18feets=18\times0.3048=5.4864m"

By definition, Total distance travelled is the sum of Thinking distance and skids or braking distance

Thus

Total distance

"D=v\times t+d"

"D=20.11\times0.5+5.4864=15.54m"